# captionUnknown document class (or package)

### v1

@@ 3 Tasks and Metrics
To understand the role of language in 3D understanding, it is important to first understand the tasks that 3D vision-language models attempt to solve. Research has grown to include a broad spectrum of research tasks, each with their own set of commonly used datasets and evaluation metrics. Here, we aim to list current 3D vision-language tasks, and their corresponding evaluation metrics. We broadly categorize the tasks by their input and output modalities.
We then begin our analysis of methods for solving these tasks in Sec. 4 and Sec. 5. Then, in Sec. 6, we detail the datasets that are currently used for training and evaluation for these tasks.
@@
Evaluation Metrics for 3D captioning require comparing the generated captions against ground truth captions for testing samples.
Exact Match (EM) requires that the generated caption exactly matches the ground truth. Exact Match has different accuracy thresholds, denoted as EM@$K$, which means that the correct answer is within the top “$K$” answers generated by the model. Commonly used thresholds are EM@1 and EM@10. However, natural language captions with the same semantic meaning can be represented in many ways, so the dominant metrics for captioning are automated text generation metrics [219] that aim to measure matching n-grams or semantic similarity rather than complete sentence matches. BLEU [220] matches n-grams between the predicted and true captions, with “BLEU@$x$” referring to matching n-grams of length “$x$” (typical values are in the range 1-4). This still requires matching exact words, but is slightly more robust to rearrangements in phrasing. ROUGE [221] similarly aims to match n-grams, with the commonly used ROUGE-L focusing on the structural similarity of sentences. METEOR [222], is based upon the precision and recall of unigram matches, with “matches” also existing between synonyms and words which are morphological variants of each other. CIDEr [223] weights n-grams by their frequency, with higher frequency n-grams given lower weights. As the above metrics rely on n-gram matches, they cannot account for different by semantically similar words. Thus, various metrics that measures semantic content overlap via similarity in learned embedding spaces (e.g. , SentenceSim [224] and BERT Score [225]) have been introduced.
For dense captioning, where captions are localized to parts of the scene, adjusted benchmarks are needed. Often, BLEU, ROUGE, METEOR and CIDEr scores are still used, however the score is set to zero if the intersection over union (IoU) between the predicted bounding box and the object is less than a threshold “$k$”. Typical “$k$” values are 0.25 and 0.5 [19, 217, 226]. However, these metrics focus on the captioning recall while ignoring false positives. This is addressed by more recent work that additionally measures the precision and F-1 score of generated captions w.r.t. the BLEU, ROUGE, METEOR and CIDEr metrics [227].
@@
Evaluation Metrics for 3D grounding require comparing predicted locations, most often in the form of bounding boxes, with ground truth locations of objects from testing samples. Acc@$K$IoU [218] is the widely used metric in 3D visual grounding, which measures the percentage of positive predictions which have a intersection over union (IoU) with the ground truth greater than a threshold $K$, which is usually set to 0.25 or 0.5. It is worth noting that some datasets evaluate the performance in different scenarios. For example, ScanRefer [218] divides the datasets into unique/multiple/overall splits. Some methods measure the average IoU [153, 232] whereas other methods measure the average distance between the centers of the bounding boxes [153]. For multi-object grounding, the F1 score is used as the metric [230]. They first get the one-to-one matching between the predicted and ground truth bounding boxes according to IoUs. Then the pairs with IoUs higher than a threshold are regarded as true positives.
@@
3D Situated Question Answering (3D-SQA) is a special case of 3D-QA. The key difference is that 3D-QA requires the model to answer questions from the perspective of a spectator with access to all information about the scene, while 3D-SQA needs the answer from the perspective of a player in a pre-defined situation. For example, 3D-SQA may ask “how many chairs are in front of me?” given the situation of “standing behind and facing the dining table”.
@@
For 3D Navigation, there are two primary metrics to evaluate the performance. 1) Success Rate (SR) measures whether the 3D agent reaches the target locations within a predefined distance threshold. 2) Success Rate Weighted by Path Length (SPL) [238], which is calculated as the SR weighted by the ratio of the ground truth length and actual path length, aims to reflect how efficiently the model achieved its goal. Other metrics include Oracle Success Rate (OSR), Trajectory Length (TL) and Goal Process (GP) [239]. In addition to the above metrics that measure whether the agent successfully reached the target and its efficiency, it is also useful to consider how well the agent path matched the path specified by the language (when language is used to specify a detailed paths). One such metric is Success weighted by normalized Dynamic Time Warping (SDTW) [240] which combines the SR with the difference between the agent path and a ground truth path for a given instruction. Please note that our discussion focuses exclusively on metrics used within 3D-LLMs methods. We encourage readers to refer to Gu et al. [241] for summary of navigation metrics.
@@ 3.5 Text-to-3D Generation (Text $\rightarrow$ 3D)
3D scene understanding tasks have been widely studied. At its core, scene understanding entails recognizing and categorizing all objects present within a designated 3D environment, a process known as semantic [246, 247, 248, 46, 249, 250] or instance-level [251, 252, 253, 254, 255, 256, 257] understanding. This stage is critical, as it forms the basis upon which more nuanced interpretations are built. Subsequently, a higher level of scene understanding focuses on spatial understanding, which refers to the construction of spatial scene graphs  [258, 259] and the semantics of object relationships [260, 261]. Going further, potential interactions can be predicted, such as affordances [262, 13, 14, 15, 263], scene changes [264, 265], and understanding the broader context of the scene, e.g. the functionality and aesthetic styles [266]. 3D data also presents unique challenges that are not present in 2D, such as the relatively high cost of obtaining and labeling 3D data, sparse 3D data structures that are not uniformly dense or aligned to a grid, and the need to reconcile multiple (possibly occluded) viewpoints of the same objects [261, 256]. To this end, researchers have leveraged the power of language, where the semantics and relationships within a 3D world can be embedded. Recent efforts in integrating Large Language Models (LLMs) with 3D data have shown promise in achieving multi-level understanding and interaction, leveraging LLMs’ inherent strengths, namely zero-shot learning, in-context learning, step-by-step reasoning, and extensive world knowledge.
@@ 4.1 How do LLMs process 3D scene information?
Given the diversity of 3D representations, as described in Sec. 2.1, there are multiple ways to obtain 3D features. As shown in 3D Geometry column in Tab. I, Point Clouds [21, 24, 19, 171, 172, 267, 268, 269, 153, 270, 271, 266, 272] are most common due to their simplicity and compatibility with various pre-trained 3D encoders, which makes them a popular choice for multi-task and multi-modal learning methods. Multi-view Images [11, 14, 17, 20, 270, 273] are also frequently used as research on 2D feature extraction is well-established, which means that 3D feature extraction requires only an additional 2D to 3D lifting scheme. RGB-D data, readily obtained using depth cameras, is commonly used in 3D embodied agent systems to extract viewpoint-relevant information for navigation and understanding. 3D Scene Graphs are a more abstract 3D representation, excelling in modeling the presence of objects and their relationship, and capturing high-level information of scenes. They are often utilized in 3D scene classification [274] and planning tasks [275]. NeRFs are currently less used in 3D-LLM methods [21]. We believe this is due to their implicit nature which makes them more difficult to tokenize and integrate with feed-forward neural networks.
@@
Current methods use different architectures (see Fig. 2) and modules that align 3D features with LLM input spaces (see 3D+LLM column in Tab. I). For models which only accept 3D inputs (Fig. 2a), a linear layer [269, 24, 266] or an MLP [171, 172] is used as the alignment module to transform 3D features into the LLM input space. Models which accept 3D and text as input often use two separate branches to align 3D features and text (Fig. 2b). Some works [171, 172] employ a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment. Others, such as [270, 271], create transformer-based alignment modules, where the standard transformer architecture is tweaked to better suit the different types of 3D data, such as dense point clouds and sparse LiDAR scans. Meanwhile, text is encoded using the pre-existing LLM text embedding table. Other works [153, 268, 276] follow the Q-Former style approach of [195] to align 3D features and text (Fig. 2c), introducing fixed length query tokens as additional input, and following a BERT-based structure to facilitate interaction between 3D and text features during alignment. Mostly, these three types of architectures described above achieve alignment by utilizing 3D captioning datasets [218], where the captioning loss, i.e.the cross-entropy loss between captions generated by LLMs and the brief, ground truth description of the scene, is used to fine-tune the alignment module, while freezing the pre-trained 3D feature extractor and LLM.
@@
Finally, some models [21, 20, 22, 23, 273, 12, 275, 277, 14, 17] employ closed-sourced models like ChatGPT and do not train an alignment module at all (Fig. 2d). Instead of aligning 3D features with the LLM input space, text descriptions are directly generated from the 3D data, such as by describing 3D bounding boxes, positions and relationships, or by using pre-existing captions. These text descriptions are input into ChatGPT. No additional alignment module is proposed in these works, and thus no training is required.

### v2

@@ 3 Tasks and Metrics
To understand the role of language in 3D understanding, it is important to first understand the tasks that 3D vision-language models attempt to solve. Research has grown to include a broad spectrum of research tasks, each with their own set of commonly used datasets and evaluation metrics. Here, we aim to summarize current 3D vision-language tasks, and their corresponding evaluation metrics. We broadly categorize the tasks by their input and output modalities, and provide an example for each category in Fig. 2. We then describe methods for solving these tasks in Sec. 4 and Sec. 5. Then, in Sec. 6, we detail the datasets that are currently used for training and evaluation for these tasks.
@@
Evaluation Metrics for 3D captioning require comparing the generated captions against ground truth captions for testing samples. Exact Match (EM) requires that the generated caption exactly matches the ground truth. Exact Match has different accuracy thresholds, denoted as EM@$K$, which means that the correct answer is within the top “KK” answers generated by the model. Commonly used thresholds are EM@1 and EM@10. However, natural language captions with the same semantic meaning can be represented in many ways, so the dominant metrics for captioning are automated text generation metrics [102] that aim to measure matching n-grams or semantic similarity rather than complete sentence matches. BLEU [103] matches n-grams between the predicted and true captions, with “BLEU@xx” referring to matching n-grams of length “xx” (typical values are in the range 1-4). This still requires matching exact words, but is slightly more robust to rearrangements in phrasing. ROUGE [104] similarly aims to match n-grams, with the commonly used ROUGE-L focusing on the structural similarity of sentences. METEOR [105], is based upon the precision and recall of unigram matches, with “matches” also existing between synonyms and words which are morphological variants of each other. CIDEr [106] weights n-grams by their frequency, with higher frequency n-grams given lower weights. As the above metrics rely on n-gram matches, they cannot account for different but semantically similar words. Thus, various metrics that measures semantic content overlap via similarity in learned embedding spaces (e.g. , SentenceSim [107] and BERTScore [108]) have been introduced.
For dense captioning, where captions are localized to parts of the scene, adjusted benchmarks are needed. Often, BLEU, ROUGE, METEOR and CIDEr scores are still used, however the score is set to zero if the intersection over union (IoU) between the predicted bounding box and the object is less than a threshold “kk”. Typical “kk” values are 0.25 and 0.5 [19, 100, 109]. However, these metrics focus on the captioning recall while ignoring false positives. This is addressed by more recent work that additionally measures the precision and F-1 score of generated captions w.r.t. the BLEU, ROUGE, METEOR and CIDEr metrics [110].
@@
Evaluation Metrics for 3D grounding require comparing predicted locations, most often in the form of bounding boxes, with ground truth locations of objects from testing samples. Acc@KKIoU [101] is the widely used metric in 3D visual grounding, which measures the percentage of positive predictions which have a intersection over union (IoU) with the ground truth greater than a threshold $K$, which is usually set to 0.25 or 0.5. It is worth noting that some datasets evaluate the performance in different scenarios. For example, ScanRefer [101] divides the datasets into unique/multiple/overall splits. Some methods measure the average IoU [115, 116] whereas other methods measure the average distance between the centers of the bounding boxes [115]. For multi-object grounding, the F1 score is used as the metric [113]. They first get the one-to-one matching between the predicted and ground truth bounding boxes according to IoUs. Then the pairs with IoUs higher than a threshold are regarded as true positives.
@@
3D Situated Question Answering (3D-SQA) is a special case of 3D-QA. The key difference is that 3D-QA requires the model to answer questions from the perspective of a spectator with access to all information about the scene, while 3D-SQA needs the answer from the perspective of a player in a pre-defined situation. For example, 3D-SQA may ask “how many chairs are in front of me?” given the situation of “facing the dining table”.
@@
For 3D Navigation, there are two primary metrics to evaluate the performance. 1) Success Rate (SR) measures whether the 3D agent reaches the target locations within a predefined distance threshold. 2) Success Rate Weighted by Path Length (SPL) [120], which is calculated as the SR weighted by the ratio of the ground truth length and actual path length, aims to reflect how efficiently the model achieved its goal. Other metrics include Oracle Success Rate (OSR), Trajectory Length (TL) and Goal Process (GP) [121]. Please note that our discussion focuses exclusively on metrics used within 3D-LLMs methods. We encourage readers to refer to Gu et al. [122] for a summary of navigation metrics.
@@ 3.5 Text-to-3D Generation (Text $\rightarrow$ 3D)
3D scene understanding tasks have been widely studied. At its core, scene understanding entails recognizing and categorizing all objects present within a designated 3D environment, a process known as semantic [127, 128, 129, 35, 130, 131] or instance-level [132, 133, 134, 135, 136, 137, 138] understanding. This stage is critical, as it forms the basis upon which more nuanced interpretations are built. Subsequently, a higher level of scene understanding focuses on spatial understanding, which refers to the construction of spatial scene graphs [139, 140] and the semantics of object relationships [141, 142]. Going further, potential interactions can be predicted, such as affordances [13, 14, 15, 143, 144], scene changes [145, 146], and understanding the broader context of the scene, e.g. the functionality and aesthetic styles [147]. 3D data also presents unique challenges that are not present in 2D, such as the relatively high cost of obtaining and labeling 3D data, working with sparse 3D data structures that are not uniformly dense or aligned to a grid, and the need to reconcile multiple (possibly occluded) viewpoints of the same objects [142, 137]. To this end, researchers have leveraged the power of language, where the semantics and relationships within a 3D world can be embedded. Recent efforts in integrating LLMs with 3D data have shown promise in achieving multi-level understanding and interaction, leveraging LLMs’ inherent strengths, namely zero-shot learning, in-context learning, step-by-step reasoning, and extensive world knowledge.
@@
3D Component LLM Component 3D+LLM Method Subsection 3D Geometry Vision Model Tuning LLM Abilities Fine-tuning LLM Base # Parameters Hardware Alignment Module Date Chen et al. [148] 4.2 SG CLIP F WK None GPT2 1.5B 1 3080 - 09/22 ConceptFusion [18] 4.2/4.4 RGB-D OpenSeg F IF/R None GPT3 175B 1 3090 - 02/23 ViewRefer [20] 4.2 MVI Multi-View Transformer T WK None GPT3 175B 4 A100 Transformer 03/23 LLM-Grounder [21] 4.2 PC/NeRF OpenScene/LERF F IF/ICL/R None GPT3.5/4 - - - 09/23 Abdelreheem et al. [22] 4.2 Mesh Vision Transformer F IF/WK None GPT3.5 - 1 3090 - 09/23 Transcribe3D [23] 4.2 PC Groupfree Transformer F IF/R None GPT3.5/4 - - - 10/23 zero-shot 3DVG [16] 4.2 RGB-D Mask3d F IF/ICL/R None GPT3.5/4 - - Transformer 11/23 3DAP [149] 4.2 MVI - - IF/ICL/R PT GPT4V - - - 12/23 SpatialPIN [150] 4.2 RGB-D LanSAM/Depth Anything F IF/R PT GPT4V - - - 03/24 SIG3D [151] 4.2 PC OpenScene F IF/ICL/R LF Flan-T5 2.7B 1 A100 MLP 10/24 3D-LLM [115] 4.3 PC/VG Mask2Former/SAM F IF/ICL/R LF OPT/Flan-T5 9B/2.7B/3B 64 V100 QFormer 07/23 Chat-3D [80] 4.3 PC Point-BERT F IF/ICL/R AF Vicuna 7B - Linear layer 08/23 LEO [152] 4.3 PC/MVI OpenClip/PointNet++ F IF/ICL/R LoRA Vicuna 7B 8 A100 Transformer 11/23 LL3DA [19] 4.3 PC/3D-BB Vote2Cap-DETR F IF/ICL/R AF OPT 1.3B 8 3090 QFormer 11/23 Point-LLM [153] 4.3 PC Point-BERT F IF/ICL/R Full LLaMA 7B/13B 8 A100 Linear layer 12/23 GPT4Point [154] 4.3 PC Point-BERT F IF/ICL/R LF OPT/Flan-T5 6.7B/3B 8 A100 QFormer 12/23 Chat-3D v2 [81] 4.3 PC Uni3D F IF/ICL/R AF Vicuna 7B 4 A40 MLP 12/23 LiDAR-LLM [155] 4.3 PC/VG VoxelNet F IF/ICL/R AF LLaMA 7B 4 A100 Transformer 12/23 Chat-Scene [156] 4.3 PC/MVI Uni-3D F IF/ICL/R LoRA Vicuna-v1.5 7B 4 A100 MLP 12/23 3DMIT [157] 4.3 PC EPCL/Uni3D F IF/ICL/R LoRA Vicuna 7B 8 A100 Linear layer 01/24 ShapeLLM [158] 4.3 PC ReCon++ F IF/ICL/R Full Vicuna 7B 8 A800 Linear layer 02/24 Scene-LLM [147] 4.3 PC/VG ConceptFusion F IF/ICL/R LF LLaMA-2 7B 32 A100 Linear layer 03/24 MiniGPT-3D [159] 4.3 PC Point-BERT F IF/ICL/R LoRA Phi-2 2.7B 1 3090 QFormer 05/24 LLaNA [160] 4.3 NeRF nf2vec F IF/ICL/R AF LLaMA-2 7B 4 A100 MLP 06/24 GreenPLM [161] 4.3 PC Uni3D F IF/ICL/R LoRA Phi-2 2.7B 1 3090 MLP 09/24 LLaVA-3D [162] 4.3 MVI/RGB-D SigLIP F IF/ICL/R Full Qwen2 7B 16 A100 MLP 09/24 Robin-3D [163] 4.3 PC/MVI Uni-3D F IF/ICL/R LoRA Vicuna-v1.5 7B 8 A6000 MLP 09/24 PerLA [164] 4.3 PC/SG Vote2Cap-DETR F IF/ICL/R AF OPT 1.3B 2 H100 QFormer 11/24 Video-3D LLM [165] 4.3 MVI/RGB-D SigLIP F IF/ICL/R Full Qwen2 7B 8 A100 Transformer 12/24 GPT4Scene [166] 4.3 MVI SigLIP F IF/ICL/R Full Qwen2 7B 8 A100 Transformer 01/25 LSceneLLM [167] 4.3 PC OpenScene F IF/ICL/R Full LLaMA-2 7B 64 V100 MLP 02/25 Inst3D-LMM [168] 4.3 PC/RGB-D Uni-3D F IF/ICL/R LoRA Vicuna-v1.5 7B 8 A100 MLP 03/25 SplatTalk [169] 4.3 MVI/3DGS SigLIP F IF/ICL/R LoRA Qwen2 7B 1 H100 MLP 03/25 Ross3D [170] 4.3 MVI SigLIP F IF/ICL/R Full Qwen2 7B 8 A100 MLP 04/25 3D-LLaVA [171] 4.3 PC 3D U-Net F IF/ICL/R LoRA Vicuna-v1.5 7B 8 3090 MLP 04/25 Spatial-MLLM [172] 4.3 MVI VGGT F IF/ICL/R Full Qwen2.5 3B 4 A800 MLP 05/25 VG LLM [173] 4.3 MVI VGGT F IF/ICL/R Full Qwen2.5 3B 8 H800 Linear layer 05/25 VLM-3R [174] 4.3 MVI CUT3R F IF/ICL/R LoRA Qwen2 7B 16 H200 MLP 05/25 LEO-VL [175] 4.3 MVI/RGB-D Qwen2.5-VL ViT F IF/ICL/R LoRA Qwen2.5 7B 8 A100 MLP 06/25 3DRS [176] 4.3 MVI/RGB-D VGGT F IF/ICL/R Full Qwen2 7B 8 H100 Transformer 06/25 Point-Bind [177] 4.4 PC I2P-MAE F IF/ICL/R LF LLaMA 7B 8 A100 Linear layer 09/23 JM3D-LLM [178] 4.4 PC PointNet++/Point-BERT F IF/ICL/R LF Vicuna 7B 3 A100 MLP 10/23 MultiPLY [24] 4.3/4.4/4.5 PC ConceptGraph T IF/ICL/R AF Vicuna 13B 128 V100 Linear layer 01/24 3DLLM-Mem [179] 4.4 MVI/RGB-D SigLIP F IF/ICL/R Full Qwen2 7B 8 TPU v5p MLP 05/25 LLM-Planner [12] 4.5 VG Seg&Depth F IF/ICL/R None GPT3 175B - - 03/23 SayPlan [180] 4.5 SG SG generator F IF/ICL/R None GPT4 - - - 07/23 VoxPoser [13] 4.5 RGB-D OWL-ViT/SAM F IF/ICL/R None GPT4 - - - 07/23 UniHSI [181] 4.5 RGB-D - - IF/ICL/R None GPT3.5/4 - 1 A100 - 09/23 LAN-grasp [14] 4.5 MVI OWL-ViT F IF/ICL/R None GPT4 - - - 10/23 Agent3D-Zero [17] 4.3/4.5 MVI - - IF/ICL/R PT GPT4V - - - 03/24 NaviLLM [11] 4.5 MVI EVA-CLIP-Large F IF/ICL/R LF Vicuna 7B 8 A100 Transformer 12/23 ManipLLM [15] 4.5 RGB-D CLIP F IF/ICL/R AF LLaMA 7B 1 A100 Linear layer 03/24 3D-VLA [182] 4.3/4.5 PC/MVI Mask2Former/SAM F IF/ICL/R LoRA Flan-T5 3B 384 V100 QFormer 03/24 PolyGen [183] 4.6 n-gon Mesh Custom Transformer F - Full Custom - 4 V100 - 02/20 LLMR [184] 4.6 Unity, Mesh Dall-E-2, CLIP F IF/ICL/R None GPT4 - 1 3080 - 09/23 3D-GPT [26] 4.6 Blender, Mesh - - IF/R None GPT3.5/4 - - - 10/23 MeshGPT [185] 4.6 Mesh GPT-2-style Transformer F - Full GPT-2-style $\sim$345M 4 A100 - 11/23 ShapeGPT [186] 4.6 SDF 3D VQ-VAE, CLIP F IF/R Full T5 - 4 A100 T5 11/23 Holodeck [187] 4.6 Mesh - - IF/R None GPT4 - 8 RTX 8k CLIP/SBERT 12/23 GALA-3D [25] 4.6 3DGS MVDream, Control-Net F IF/R None Any (GPT3.5) - - - 02/24 LLaMA-Mesh [188] 4.6 Mesh - F IF/R Full LLaMA-3.1 8B 32 A100 MLP 11/24
TABLE I: Summarization of 3D-LLMs methods. The 3D Geometry column collects the 3D geometric information being used for each method, such as Point Cloud (PC), Multi-view Images (MVI), RGB+Depth (RGB-D), 3D Bounding Box (3D-$B$), Scene Graph (SG), Voxel Grid (VG), Gaussian Splatting (3DGS) and NeRF. The tuning column outlines whether the vision model is finetuned during training (True/False). IF, ICL, R, and WK denote the LLM Abilities of Instruction-following, In-Context Learning, Reasoning, and World Knowledge. The Fine-tuning column summarizes how LLM components are fine-tuned, such as Prompt Tuning (PT), Low-Rank Adaptation (LoRA), Adaptive Fine-tuning (AF), Layer Freezing (LF), or Full Fine-tuning (Full). The Hardware column shows methods that are trained either with numbers of Nvidia GPU and the specific GPU type or with no training involved.
@@ 4.1 How do LLMs process 3D scene information?
Given the diversity of 3D representations, as described in Sec. 2.1, there are multiple ways to obtain 3D features. As shown in 3D Geometry column in Tab. I, Point Clouds [21, 24, 19, 80, 81, 153, 154, 157, 115, 152, 155, 147, 177] are most common due to their simplicity and compatibility with various pre-trained 3D encoders, which makes them a popular choice for multi-task and multi-modal learning methods. Multi-view Images [11, 14, 17, 20, 152, 149] are also frequently used as research on 2D feature extraction is well-established, which means that 3D feature extraction requires only an additional 2D to 3D lifting scheme. RGB-D data, readily obtained using depth cameras, is commonly used in 3D embodied agent systems to extract viewpoint-relevant information for navigation and understanding [162, 165, 168, 175, 176]. 3D Scene Graphs provide more abstract 3D representation, excelling in modeling the presence of objects and their relationship, and capturing high-level information of scenes. They are often utilized in 3D scene classification [148], understanding [164] and planning tasks [180]. NeRFs are currently less used in 3D-LLM methods [21, 160]. We believe this is due to their implicit nature which makes them more difficult to tokenize and integrate with feed-forward neural networks.
@@
Current methods use different architectures (see Fig. 3) and modules that align 3D features with LLM input spaces (see 3D+LLM column in Tab. I). For models which only accept 3D inputs (Fig. 3a), a linear layer [157, 24, 147] or an MLP [80, 81] is used as the alignment module to transform 3D features into the LLM input space. Models which accept 3D and text as input often use two separate branches to align 3D features and text (Fig. 3b). Some works [80, 81] employ a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment. Others, such as [152, 155], create transformer-based alignment modules, where the standard transformer architecture is tweaked to better suit the different types of 3D data, such as dense point clouds and sparse LiDAR scans. Meanwhile, text is encoded using the pre-existing LLM text embedding table. Other works [115, 154, 182] follow the Q-Former style approach of [87] to align 3D features and text (Fig. 3c), introducing fixed length learnable query tokens as additional input, and following a BERT-based structure to facilitate interaction between 3D and text features during alignment. Mostly, these three types of architectures described above achieve alignment by utilizing 3D captioning datasets [101], where the captioning loss, i.e., the cross-entropy loss between captions generated by LLMs and the brief, ground truth description of the scene, is used to fine-tune the alignment module, while freezing the pre-trained 3D feature extractor and LLM.
@@
Finally, some models [21, 20, 22, 23, 149, 12, 180, 181, 14, 17] employ closed-source models like ChatGPT and do not train an alignment module at all (Fig. 3d). Instead of aligning 3D features with the LLM input space, text descriptions are directly generated from the 3D data, such as by describing 3D bounding boxes, positions and relationships, or by using pre-existing captions. These text descriptions are input into ChatGPT. No additional alignment module is proposed in these works, and thus no training is required.